OpenStudy (anonymous):

A car depreciates (loses value) at a rate of 4.5% annually. Greg purchased a car for $12,500. Which equation can be used to determine the value of the car, V, after 5 years?

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

USe future value equation ...FT=PT(1+rt)

OpenStudy (cwrw238):

use V = 12,500(1 - r)^n where r = annual rate of depreciation and n = number of years

OpenStudy (cwrw238):

r can be written as a decimal - i e 4.5 / 100 = .045 and n = 5